You are given the following information for Wine and Cork Enterprises (WCE): rRF = 5%; rM = 9%; RPM = 4%, and beta = 1

1.What is WCE's required rate of return? Do not round intermediate calculations. Round your answer to two decimal places.

2. If inflation increases by 1% but there is no change in investors' risk aversion, what is WCE's required rate of return now? Do not round intermediate calculations. Round your answer to two decimal places.

3.Assume now that there is no change in inflation, but risk aversion increases by 1%. What is WCE's required rate of return now? Do not round intermediate calculations. Round your answer to two decimal places.

4.If inflation increases by 1% and risk aversion increases by 1%, what is WCE's required rate of return now? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

1.
=5%+1*4%=9.00%

2.
=5%+1%+1*4%=10.00%

3.
=5%+1*(4%+1%)=10.00%

4.
=5%+1%+1*(4%+1%)=11.00%